# Harnessing the Unseen: The Hidden Influence of Intrinsic Knowledge in Long-Context Language Models

### v1

@@ Abstract
Recent advances in long-context models (LCMs), designed to handle extremely long input contexts, primarily focus on utilizing external contextual information, often leaving the influence of large language models’ intrinsic knowledge underexplored. In this work, we investigate how this intrinsic knowledge affects content generation and demonstrate that its impact becomes increasingly pronounced as context length extends. Furthermore, we show that the model’s ability to utilize intrinsic knowledge, which we call intrinsic retrieval ability, does not improve simultaneously with its ability to leverage contextual knowledge through extrinsic retrieval ability. Moreover, better extrinsic retrieval can interfere with the model’s ability to use its own knowledge effectively, limiting its full potential. To bridge this gap, we design a simple yet effective Hybrid Needle-in-a-Haystack test that evaluates models based on their capabilities across both retrieval abilities, rather than solely emphasizing extrinsic retrieval ability. Our experimental results reveal that Qwen-2.5 models significantly outperform Llama-3.1 models, demonstrating superior intrinsic retrieval ability. Moreover, even the more powerful Llama-3.1-70B-Instruct model fails to exhibit better performance under LCM conditions, highlighting the importance of evaluating models from a dual-retrieval perspective.
1 Introduction
@@
Recent advancements in open-source models such as LLaMA (Grattafiori et al., 2024), Qwen (Yang et al., 2024a; b) and closed-source models such as GPT-4 (Achiam et al., 2023; OpenAI, 2024), Claude (Anthropic, 2024) and Gemini (DeepMind, 2024) have all incorporated long-context capabilities, extending their context windows to tens of thousands of tokens. For instance, GPT-4o and Claude 3.5 have expanded their context windows to 128K and 200K tokens respectively. The emergence of those Long Context Models (LCMs) has significantly advanced the ability of language models to process and generate coherent content over extended inputs. These models have become increasingly useful for various tasks, including document summarization (Chang et al., 2023; Yen et al., 2025), multi-turn conversations (Maharana et al., 2024) and question answering (Yang et al., 2018; An et al., 2023; Karpinska et al., 2024), where retaining and utilizing long-term dependencies is essential.
@@
Numerous benchmarks have been proposed to evaluate LCMs’ effectiveness in handling long contexts. For example, the Needle-in-a-Haystack test (Kamradt, 2023) inserts critical information (i.e. “needle”) in long irrelevant documents (i.e. “haystack”) and test LCMs’ ability to retrieve the information. Building upon this experiment, benchmarks such as RULER (Hsieh et al., 2024), InfiniteBench (Zhang et al., 2024) and HELMET (Yen et al., 2025) extend to more realistic tasks such as QA, summarization etc., and very long contexts (100K+ tokens). These benchmarks systematically assess LCMs’ capacity to process and generate meaningful outputs over extraordinarily long contexts.
However, the aforementioned benchmarks for LCMs primarily evaluate how well these models leverage external long-context inputs, without considering the influence of their parametric (intrinsic) knowledge (Roberts et al., 2020; Wang et al., 2021; Jiang et al., 2020). Parametric knowledge has been shown to be critical in language models for short context tasks (Xie et al., 2023; Xu et al., 2024b) and retrieval-augmented generation (Mallen et al., 2023; Wang et al., 2024a). In this work, we argue that intrinsic knowledge plays a crucial role in long-context generation and can significantly impact model performance. Concretely, we create a new dataset - IWhoQA - based on WhoQA (Pham et al., 2024) and test the behavior of LCMs when their intrinsic knowledge either supports or conflicts with the extrinic knowledge in the long context input. We show that LCMs consistently achieve better performance when the context supports their intrinsic knowledge and they gradually rely more on intrinsic knowledge when the input context length is increased.
Given the importance of intrinsic knowledge for LCMs, we further seek to determine whether enhancing a model’s extrinsic retrieval ability can influence the expression of intrinsic knowledge. Specifically, we compare between RoPE (Su et al., 2021) and STRING (An et al., 2024b), a modified version of RoPE designed for long-context tasks, on a curated dataset collected from HotPotQA (Yang et al., 2018). We show that STRING improves extrinsic retrieval ability but unintentionally suppresses intrinsic retrieval ability, even when the latter positively contributes to generation quality. This finding highlights a trade-off between utilizing external long-context inputs and leveraging the LCMs’ intrinsic knowledge.
@@
Based on these observations, we propose a simple yet efficient Hybrid Needle-in-a-Haystack test to offer a comprehensive assessment of how well models can integrate the intrinsic and extrinsic retrieval ability during generation. Specifically, we design a set of questions such as “What’s the favorite thing of the person who wrote {Book_Name}?” which require the model to first retrieve the author’s name based on intrinsic knowledge and then retrieve a needle to answer the question. Our experimental results show that the ability of the Qwen2.5 family to integrate both intrinsic and extrinsic knowledge scales linearly with model size. In contrast, the Llama3.1 family fails to achieve noticeable improvement even with a substantial increase in model parameters, indicating that Llama3.1 struggles to effectively leverage its intrinsic knowledge in long-context scenarios. Our contributions can be summarized as follows:
@@
  • We present the first study demonstrating that the intrinsic knowledge of LCMs plays a significant role in long-context generation, with its influence becoming increasingly pronounced as the context length grows.
@@
  • We show that naively improving extrinsic retrieval ability will unintentionally suppress intrinsic retrieval ability, revealing a trade-off between utilizing external inputs and leveraging intrinsic knowledge in LCMs.
@@
  • We introduce a novel Hybrid Needle-in-a-Haystack test to examine LCMs’ ability to integrate the intrinsic and extrinsic retrieval ability during generation. We show that this ability doesn’t always scale with the model size, highlighting limitations in existing LCMs.
2 Related Work
@@
Since training transformers on long context from scratch is prohibitively expensive, research has focused on extending the context window of existing transformers through two broad categories: 1) Finetuning on a small amount of long context data (Peng et al., 2024; Lazarevich et al., 2025; Chen et al., 2023; jquesnelle, 2023) and 2) Training-free methods based on reusing relative positions seen during training (An et al., 2024a; bloc97, 2023; emozilla, 2023; Su, 2023; Jin et al., 2024). Self-Extend (Jin et al., 2024), Dual Chunk Attention (An et al., 2024a), and STRING (An et al., 2024b) all propose replacing large relative distances between query and key tokens by overwriting them with proportionately smaller relative distances seen during training. Relative distances below a certain threshold remain unmodified.
Another challenge for long context language modeling is storing key and value activations for all past tokens (the KV cache). Significant research has focused on selectively attending to only parts of the context with sparse attention (Han et al., 2024; Xiao et al., 2024; Jiang et al., 2024). Additional optimization comes from estimating the importance of the KV cache and evicting low-importance cache entries (Fu et al., 2024; Feng et al., 2025; Cai et al., 2024; Li et al., 2024).
@@
Multiple synthetic tasks have been proposed to evaluate LCMs, including Needle-in-a-Haystack (NIAH) test (Kamradt, 2023), long-context QA, summarization, reasoning, extrinsic learning (ICL) and retrieval-augmented generation (RAG), among others (Hsieh et al., 2024; Kim et al., 2025; Gao et al., 2025; Liu et al., 2024; Yen et al., 2025). Variants of NIAH include multiple-needle, long-needle, and needle-in-needle configurations (Gao et al., 2025) and have also been extended to test long-context reasoning (Kuratov et al., 2024), mathematical (Wang et al., 2024b) and multilingual capabilities (Hengle et al., 2024). However, several works have found that the standard NIAH task may not truly reflect an LCM’s downstream capabilities and have proposed comprehensive benchmarks such as RULER (Hsieh et al., 2024) and it’s multilingual extension ONERULER (Kim et al., 2025), LongBench (Bai et al., 2024) , LongGenBench (Liu et al., 2024), L-Eval (An et al., 2023), InfiniteBench (Zhang et al., 2024) and HELMET (Yen et al., 2025). These benchmarks contain several tasks such as citation generation, reranking, QA, summarization, ICL, and RAG. Karpinska et al. (2024) introduced NoCha, a dataset of human-written true and false claims about recently published fictional books. While previous work has primarily studied long context in isolation, some going so far as to use entirely fictional contexts (Gao et al., 2025; Karpinska et al., 2024), to the best of our knowledge the problem of how an LLM’s intrinsic knowledge affects its long context abilities has not been explored.
Intrinsic Knowledge
During pre-training, LLMs memorize a vast amount of knowledge into their parameters, known as parametric (intrinsic) knowledge (Roberts et al., 2020; Wang et al., 2021; Jiang et al., 2020). This parametric knowledge can often conflict with new information presented via prompting through a phenomenon known as knowledge conflict (Xie et al., 2023; Longpre et al., 2022; Pan et al., 2023; Xu et al., 2024a). While earlier works have found that LLMs rely on parametric knowledge when prompted with simple counter-factual statements (Longpre et al., 2022; Chen et al., 2022), more recent research have demonstrated that LLMs rely more on the information presented as the context (Xie et al., 2023; Pan et al., 2023; Wan et al., 2024). Pan et al. (2023) uncovers LLMs’ sensitivity to small amounts of even misinformation in context and Xu et al. (2024a) applies prompting to persuade LLMs to disregard their parametric knowledge. However, Xie et al. (2023) further finds that when presented with conflicted evidence, LLMs resolve in favor of parametric knowledge. Zhou et al. (2023) proposes framing the conflicting information as user opinion to prevent the LLM from taking it as fact. We refer the reader to Xu et al. (2024b) for further details. However, to the best of our knowledge, the relation between parametric knowledge and long-context tasks has not been studied.
3 The Role of Intrinsic Knowledge in Long-Context Generation
Current benchmarks for LCMs focus primarily on assessing a model’s ability to extract important information or perform reasoning based only on the provided context. In contrast, we seek to investigate the role of the LCMs’ intrinsic knowledge. This section begins with the question: What role does the intrinsic knowledge of LCMs play in long-context generation? We aim to explore how the model’s intrinsic knowledge interacts with contextual data, potentially influencing the quality and accuracy of the generated results.
3.1 Dataset Creation
We design new dataset to first investigate whether LCMs naturally incorporates intrinsic knowledge during generation. Recognizing that existing datasets within the knowledge-conflict (Xie et al., 2023; Xu et al., 2024b; Pham et al., 2024) also aim to identify and investigate intrinsic knowledge of LLMs, we chose WhoQA (Pham et al., 2024), a QA dataset which contains several context-answer pairs for each question, as our initial dataset.
For each example in WhoQA, we generated multiple answers based on different phrasings of question and retained only those examples where the model generates only one stable answer. This methodology ensures that the intrinsic knowledge in the model contains only one answer for this example. Given that the intrinsic knowledge may vary among different LCMs, we collect separate datasets for each LCMs. Consequently, we constructed datasets comprising 300 examples for each model, which we refer to as the I-WhoQA dataset.
As our objective is to assess how intrinsic knowledge affects long-context generation, we select two different context-answer pairs for each example in I-WhoQA: (1) Context that aligns with the model’s intrinsic knowledge, referred to as I-WhoQA-internal subset. (2) Context that conflicts with the model’s intrinsic knowledge, referred to as I-WhoQA-conflict subset. By using these two different subsets from I-WhoQA dataset, we can uncover how intrinsic knowledge influences the generation in long-context scenarios. See Appendix 6 for an example of our I-WhoQA dataset.
@@ 3.1 Dataset Creation
Figure 1: Left) LCMs struggle to retrieve the answer from the context on the I-WhoQA-Conflict-subset i.e. when the context information conflicts with its intrinsic knowledge. Right) The upward trend of I-WhoQA-conflict (red) shows that when the intrinsic knowledge conflicts with the context, the likelihood of a LCM relying on intrinsic knowledge steadily increases with larger contexts.
3.2 Experiment
@@
To validate the role of intrinsic knowledge in long-context generation, we conducted analyses using two curated I-WhoQA subsets mentioned above, alongside two different backend models — mistral-7b-instruct-v0.2 (Jiang et al., 2023) and Llama3.1-8B-instruct (Grattafiori et al., 2024). For each backend model, we used greedy decoding strategy with a maximum of 32 generation tokens. To examine the impact of context lengths, which is key factor in long-context generation, we report experimental results with varying context lengths. For the same example, the only difference across these different context settings is the amount of irrelevant contents inserted into the context, which is consistent with the setup of Needle-in-a-Haystack test.
@@
The results in Figure 1 show that the intrinsic knowledge of LCMs will influence its generation and can be more significant with the increase of context length. Firstly, the left side of Figure 1 (Answer Matches Context) shows the performance on the IWhoQA-conflict is consistently worse than on IWhoQA-internal subset. This demonstrates that the LCMs are struggling to answer according to the context when its intrinsic knowledge contains a different answer. Secondly, from the right side of Figure 1 (Answer Conflict With Context) on the I-WhoQA-conflict subset, where conflicting contexts are provided as supporting facts, we observed a notable increase in the proportion of outputs adhering to intrinsic knowledge as context length increased. This trend is especially pronounced when Llama-3.1-8B-Instruct serve as the backend model, which can handle longer context lengths.
4 The Trade-off Between Intrinsic and Extrinsic Retrieval Abilities
In the previous section, we demonstrated that the intrinsic knowledge within LCMs plays a crucial role in long-context generation. This influence becomes increasingly significant as the context length grows. While existing research (An et al., 2024b; Han et al., 2024; emozilla, 2023) on long-context modeling has largely focused on enhancing models’ extrinsic retrieval abilities, there has been little analysis of the influence of LCMs’ intrinsic knowledge during long-context generation, which we refers to as intrinsic retrieval ability. In this section, we aims to solve the question: Will the expression of intrinsic knowledge be influenced by the extrinsic retrieval ability in long-context generation? We study the effect of STRING (An et al., 2024b), an improved position encoding method, on extrinsic and intrinsic retrieval ability of LCMs. Our experimental results reveal that simply enhancing extrinsic retrieval ability with STRING will inadvertently alter the expression of intrinsic knowledge, which in some cases can lead to a decline in model performance.
4.1 Dataset Creation
In Section 3, we built the I-WhoQA dataset, ensuring that all examples can be answered using the model’s intrinsic knowledge. Based on the I-WhoQA dataset, we further constructed the I-WhoQA-irrelevant subset. In this subset, instead of providing context relevant to the current question, we inserted completely irrelevant context to assess how models handle unrelated information and if STRING can handle the situation that need intrinsic retrieval ability rather than extrinsic retrieval ability.
The I-WhoQA-irrelevant subset represents a relatively extreme setting and may occur infrequently in real-world scenarios. More commonly, the context may contain content that is related to the question but cannot actually be used to derive the correct answer. To address this issue, we construct a corresponding dataset from HotpotQA dataset (Yang et al., 2018), where the context includes seemingly relevant but ultimately unhelpful information for answering the question. Based on the HotpotQA dataset, we construct two subsets by examining the relationship between the paragraphs provided by HotpotQA, the model’s intrinsic knowledge and the reference answer.
First, to distinguish the role of provided context from that of the model’s intrinsic knowledge, we ensured that the answers derived from the given paragraphs and the model’s intrinsic knowledge are not the same. Building on this distinction, we obtained two final subsets. The first subset, referred to as the HotpotQA-context subset, contains examples where the reference answer can be derived from the given passages rather than the intrinsic knowledge. The second subset, known as the HotpotQA-internal subset, consists of examples where the reference answer matches the model’s intrinsic knowledge. We constructed a dataset consisting of 200 examples for each subset, except for the HotpotQA-internal subset for Llama3.1-8B-Instruct, which contains a total of 93 examples.
@@ 4.1 Dataset Creation
Figure 2: I-WhoQA-irrelevant subsets: LCMs are often unable to ignore context not relevant to the question and recall the answer from their intrinsic knowledge. As the length of the context is increased, LCMs gradually start ignoring the context and generating according to their intrinsic knowledge. However, using enhanced positional encoding designed for long context tasks such as STRING causes LCMs to re-focus on irrelevant context and generate wrong answers.
Figure 3: Performance of LCMs on HotPotQA-context and -internal subsets. Left) STRING improves performance on the HotPotQA-context subset by enhancing extrinsic retrieval ability. Right) STRING hinders the model’s ability to recall intrinsic knowledge and leads to the decrease of performance.
4.2 Experiment
@@
We use the same backend models and generation configuration as described in Section 3, applied to the newly constructed datasets. For the STRING method, we configure the local value to 128 and two different the shift-ratios: 0.25 and 0.33. We report the average performance across these two settings. More details about STRING can be found in An et al. (2024b).
@@
Figure 2 shows that STRING consistently achieves lower exact match accuracy than the baseline RoPE method. We hypothesize that this degradation may result from STRING enhancing the extrinsic retrieval ability under long-context scenarios. As a result, even when the provided context is irrelevant, the STRING method will still attempts to retrieve and rely information from the context rather than using the model’s intrinsic knowledge that contains the correct answer.

### v2

@@ Abstract
Recent advances in long-context language models (LCLMs), designed to handle extremely long contexts, primarily focus on utilizing external contextual information, often leaving the influence of language models’ parametric knowledge underexplored. In this work, we firstly investigate how this parametric knowledge affects content generation and demonstrate that its impact becomes increasingly pronounced as context length extends. Furthermore, we show that the model’s ability to utilize parametric knowledge, which we call parametric recall ability, does not improve simultaneously with its ability to leverage contextual knowledge through extrinsic retrieval ability. Moreover, better extrinsic retrieval ability can interfere with the model’s parametric recall ability, limiting its full potential. To bridge this gap, we design a simple yet effective Hybrid Needle-in-a-Haystack test that evaluates models based on their capabilities across both abilities, rather than solely emphasizing extrinsic retrieval ability. Our experimental results reveal that Qwen-2.5 models significantly outperform Llama-3.1 models, demonstrating superior potential to combine various abilities. Moreover, even the more powerful Llama-3.1-70B-Instruct model fails to exhibit better performance, highlighting the importance of evaluating models from a dual-ability perspective.
Codehttps://github.com/FYYFU/Hybrid-NIAH
Introduction
@@
Recent advancements in both open-source models (e.g., LLaMA (Grattafiori et al. 2024), Qwen (Yang et al. 2024a, b)) and closed-source models (e.g., GPT-4 (Achiam et al. 2023; OpenAI 2024), Claude (Anthropic 2024), Gemini (DeepMind 2024)) have incorporated long-context capabilities and significantly extended their context windows. For instance, GPT-4o and Claude Sonnet 4 have expanded their context windows to 128K and 200K tokens respectively. The emergence of those Long Context Language Models (LCLMs) has significantly advanced the ability of language models to process and generate coherent content over extended contexts. They have become increasingly useful for various tasks, including document summarization (Yen et al. 2025), multi-turn conversations (Maharana et al. 2024) and question answering (Karpinska et al. 2024), where retaining and utilizing long-term dependencies is essential.
@@
With the advance of LCLMs, numerous benchmarks have been proposed to evaluate their effectiveness in handling long contexts. For example, the Needle-in-a-Haystack test (Kamradt 2023) inserts fabricated critical information (i.e. “needle”) into long irrelevant documents (i.e. “haystack”) to examine LCLMs’ external retrieval ability and prevent the influence of their own parametric knowledge. Other benchmarks such as RULER (Hsieh et al. 2024), InfiniteBench (Zhang et al. 2024) and HELMET (Yen et al. 2025) extend to more realistic tasks such as QA, summarization etc., and ultra long contexts (100K+ tokens) to systematically assess LCLMs’ capacity to process and generate meaningful outputs over extraordinarily external long contexts. These benchmarks evaluate how effectively these models leverage external long contexts, thereby emphasizing their extrinsic retrieval ability while overlooking the role of parametric knowledge (Roberts et al. 2020; Wang et al. 2021; Jiang et al. 2020).
However, the parametric knowledge has already been shown to be pivotal to the performance of language models, especially when parametric and extrinsic knowledge contradict each other (Xie et al. 2023; Wang et al. 2024a; Xu et al. 2024b). In this work, we argue that parametric knowledge plays an important role during long-context generation, with its impact becoming more pronounced as context length extends. To validate this hypothesis, we construct a new dataset, IWhoQA, based on WhoQA (Pham et al. 2024), specifically designed to systematically probe LCLMs in scenarios where parametric knowledge either aligns with or contradicts the external context. Experimental results show that LCLMs consistently perform better when the external context aligns with their parametric knowledge. The performance gap increases with longer contexts, reaching up to 10 points on Llama-3.1-8B-Instruct, suggesting a growing reliance on parametric knowledge during long-context generation.
Given the importance of parametric knowledge for LCLMs and the lack of its evaluation in existing benchmarks, it remains unclear whether methods that perform well in these benchmarks improve only extrinsic retrieval ability, or also enhance parametric recall ability. To address this gap, we systematically investigate the relationship between extrinsic retrieval and parametric recall. Specifically, we compare RoPE (Su et al. 2021) and STRING (An et al. 2024b), an improved variant of RoPE designed for long-context tasks, across various curated datasets featuring diverse knowledge alignments between external context and parametric knowledge. Our results show that while STRING improves extrinsic retrieval ability, it suppresses parametric recall even when parametric knowledge benefits generation. This reveals a trade-off between leveraging external context and utilizing parametric knowledge, an aspect overlooked by current benchmarks.
@@
Building on these observations, we propose a simple yet effective Hybrid Needle-in-a-Haystack test to comprehensively evaluate how well models integrate parametric recall ability with extrinsic retrieval ability during long-context generation. Specifically, we design queries such as “What’s the favorite thing of the person who wrote {Book_Name}?”—which require the model to first recall the author’s name from its parametric knowledge, and then retrieve the inserted needle in the context to answer the question. Experimental results show that the Qwen2.5 family demonstrates a near-linear improvement in its ability to integrate both types of knowledge as model size increases. In contrast, the Llama3.1 family exhibits minimal gains despite a substantial increase in parameters, suggesting that it struggles to effectively leverage parametric knowledge in long-context settings. Our contributions can be summarized as follows:
@@
  • We present the first study demonstrating that the parametric knowledge of LCLMs plays an important role in long-context generation, with its influence becoming increasingly pronounced as the context length extends.
@@
  • We show that naively improving LCLMs’ extrinsic retrieval ability will unintentionally suppress their parametric recall ability, revealing a trade-off between utilizing external contexts and leveraging parametric knowledge.
@@
  • We introduce a simple yet effective Hybrid Needle-in-a-Haystack test to examine LCLMs’ ability to integrate the parametric recall ability and extrinsic retrieval ability during long-context generation. We show that this ability doesn’t always scale with the model size, highlighting limitations in existing LCLMs.
Related Work
Parametric Knowledge
During pre-training, LLMs memorize a vast amount of knowledge into their parameters, known as parametric (intrinsic) knowledge (Roberts et al. 2020; Wang et al. 2021; Jiang et al. 2020). This parametric knowledge can often conflict with new information presented via prompting through a phenomenon known as knowledge conflict (Xie et al. 2023; Longpre et al. 2022; Pan et al. 2023; Xu et al. 2024a). While earlier works have found that LLMs rely on parametric knowledge when prompted with simple counter-factual statements (Longpre et al. 2022; Chen et al. 2022), more recent research has demonstrated that LLMs rely more on the information presented as the context (Xie et al. 2023; Pan et al. 2023; Wan et al. 2024). Pan et al. (2023) uncovers LLMs’ sensitivity to small amounts of even misinformation in context and Xu et al. (2024a) applies prompting to persuade LLMs to disregard their parametric knowledge. However, Xie et al. (2023) further finds that when presented with conflicted evidence, LLMs resolve in favor of parametric knowledge. Zhou et al. (2023) proposes framing the conflicting information as user opinion to prevent the LLM from taking it as fact. We refer the reader to Xu et al. (2024b) for further details. However, to the best of our knowledge, the relation between parametric knowledge and long-context tasks has not been studied.
@@
Training transformers on long context from scratch is prohibitively expensive. Much research has focused on extending the context window of existing models through two categories: 1) Finetuning on a small amount of long context data (Peng et al. 2024; Lazarevich et al. 2025; Chen et al. 2023; jquesnelle 2023) and 2) Training-free methods based on reusing position information. (An et al. 2024a; bloc97 2023; emozilla 2023; Su 2023; Jin et al. 2024). For example, Self-Extend (Jin et al. 2024), Dual Chunk Attention (An et al. 2024a), and STRING (An et al. 2024b) all propose replacing large relative positions by overwriting them with proportionately smaller ones seen during training. By mapping large relative positions to smaller ones that are well represented during pre-training, the model can more effectively integrate provided information into long-context generation.
@@
Multiple synthetic tasks have been proposed to evaluate LCLMs, including Needle-in-a-Haystack (NIAH) test (Kamradt 2023), long-context QA, summarization, reasoning, extrinsic learning (ICL) and retrieval-augmented generation (RAG), among others (Hsieh et al. 2024; Kim et al. 2025; Gao et al. 2025; Liu et al. 2024; Yen et al. 2025). Variants of NIAH include multiple-needle, long-needle, and needle-in-needle configurations (Gao et al. 2025) and have also been extended to test long-context reasoning (Kuratov et al. 2024), mathematical (Wang et al. 2024b) and multilingual capabilities (Hengle et al. 2024). However, several works have found that the standard NIAH task may not truly reflect an LCLM’s downstream capabilities and have proposed comprehensive benchmarks such as RULER (Hsieh et al. 2024) and its multilingual extension ONERULER (Kim et al. 2025), LongBench (Bai et al. 2024) , LongGenBench (Liu et al. 2024), L-Eval (An et al. 2023), InfiniteBench (Zhang et al. 2024) and HELMET (Yen et al. 2025). These benchmarks contain several tasks such as citation generation, reranking, QA, summarization, ICL, and RAG. Karpinska et al. (2024) introduced NoCha, a dataset of human-written true and false claims about recently published fictional books. While previous work has primarily studied long context in isolation, some going so far as to use entirely fictional contexts (Gao et al. 2025; Karpinska et al. 2024), to the best of our knowledge the problem of how an LCLM’s parametric knowledge affects its long context abilities has not been explored.
@@
Figure 1: We only count answers that align with either the injected knowledge (whether parametric or conflictual) or the opposing knowledge source. Left) LCLMs struggle to retrieve the answer from the context on the I-WhoQA-Conflict subset, i.e., when the context information conflicts with their parametric knowledge. Right) The upward trend of I-WhoQA-conflict (red) shows that when the parametric knowledge conflicts with the context, the likelihood of a LCLM relying on parametric knowledge steadily increases with larger contexts.
The Role of Parametric Knowledge in Long-Context Generation
Existing benchmarks for evaluating LCLMs primarily focus on assessing their extrinsic retrieval ability—how well they extract information or perform reasoning based solely on the external context. In contrast, our work aims to investigate the role of parametric knowledge embedded within LCLMs. We begin by posing a central question: What role does the parametric knowledge of LCLMs play in long-context generation? Specifically, we explore how parametric knowledge interacts with external context and how this interplay affects the quality of the generated output.
Dataset Creation
While existing datasets in the knowledge conflict (Xie et al. 2023; Xu et al. 2024b; Pham et al. 2024) also aim to examine the role of parametric knowledge in LLMs, they are not designed for evaluating LCLMs. To address this gap, we first construct a dataset to investigate whether LCLMs naturally incorporate parametric knowledge during long-context generation across varying context lengths. Specifically, we choose WhoQA (Pham et al. 2024), a short-form QA dataset that provides multiple context-answer pairs for each entity, as the foundation for our construction.
Specifically, for each entity in WhoQA, we generated an answer to every associated question in the dataset and retained only those entities where the model consistently produced a single, invariant answer. This filtering process ensures that the model’s parametric knowledge encodes a single, unambiguous answer for the retained entities. Since parametric knowledge may differ across LCLMs, we constructed a separate dataset of 300 examples for each model, which we refer to as the I-WhoQA dataset.
To assess how parametric knowledge influences long-context generation, we select two distinct context-answer pairs for each example in the I-WhoQA dataset: (1) a context that aligns with the model’s parametric knowledge, referred to as the I-WhoQA-Parametric subset; (2) a context that conflicts with the model’s parametric knowledge, referred to as the I-WhoQA-Conflict subset. By comparing model performance across these two subsets and varying context lengths, we aim to reveal the extent to which parametric knowledge impacts long-context generation. An example from the I-WhoQA dataset is provided in the Appendix Figure  6.
Experiment
@@
To validate the role of parametric knowledge in long-context generation, we conducted experiments using two curated I-WhoQA subsets mentioned above, alongside two different backend models — mistral-7b-instruct-v0.2 (Jiang et al. 2023) and Llama3.1-8B-instruct (Grattafiori et al. 2024). For each backend model, we employed greedy decoding strategy with a maximum of 32 generation tokens. To investigate the effect of context length, a key factor in long-context generation, we report results across varying context lengths. For each example, the only difference across these different context settings is the amount of irrelevant content inserted into the context, which is consistent with the setup of Needle-in-a-Haystack test. In our evaluation, we only include answers that clearly align with one of two competing knowledge sources: the injected context (which can represent either parametric or conflict knowledge) or the alternative source (i.e., the one not injected). This ensures a clean comparison between model behavior in context-following vs. parametric recall scenarios.
Figure 2: Performance of LCLMs on HotPotQA-Context and -Parametric subsets. Left) STRING improves performance on the HotPotQA-context subset by enhancing extrinsic retrieval ability. Right) STRING hinders the model’s ability to recall parametric knowledge and leads to the decrease of performance.
@@
Results in Figure 1 show that the parametric knowledge of LCLMs influences long-context generation and this influence becomes more pronounced as context length increases. First, the left side of Figure 1 (Answer Aligned With the Inserted Knowledge) shows the performance on the IWhoQA-Conflict subset is consistently worse than on IWhoQA-Parametric subset. This demonstrates that LCLMs are struggling to follow the given context when its parametric knowledge contains a different answer. Second, from the right side of Figure 1 (Answer Aligned With the Opposing Knowledge) on the I-WhoQA-Conflict subset, where conflicting contexts are provided as supporting facts, we observed a notable increase in the proportion of outputs adhering to opposing knowledge (i. e. parametric knowledge) as context length increased. This trend is especially pronounced when Llama-3.1-8B-Instruct serves as the backend model, which can handle longer context lengths. This suggests that as the context length increases, the model increasingly relies on its parametric knowledge even when that knowledge contradicts the external context.
The Trade-off Between Parametric Recall Ability and Extrinsic Retrieval Ability
In Section 1, we demonstrated that the parametric knowledge within LCLMs increasingly influences generation as context length grows. While existing benchmarks and research (An et al. 2024b; Han et al. 2024; emozilla 2023) on LCLMs have largely focused on evaluating and improving extrinsic retrieval, little attention has been paid to how parametric knowledge affects long-context generation. In this section, we aim to answer the question: Can we improve extrinsic retrieval without compromising parametric recall in long-context generation? We examine whether STRING (An et al. 2024b), an improved variant of RoPE designed for long-context modeling, can enhance both capabilities. Experimental results reveal that although STRING improves extrinsic retrieval, it can inadvertently impair parametric recall, sometimes leading to a decline in overall model performance.
@@
Figure 3: I-WhoQA-Irrelevant subsets: As the length of the context is increased, LCLMs ignore the context and generate according to their parametric knowledge. However, using STRING causes LCLMs to re-focus on irrelevant context and generate wrong answers.
Figure 4: Upper) Needle-in-a-Haystack. It involves directly inserting the answer into the haystack and retrieving it. Lower) Hybrid Needle-in-a-Haystack. It requires a two-step process: first, performing parametric recall to identify the retrieval target based on the model’s parametric knowledge, and then retrieving the answer from the haystack.
Dataset Creation
In Section 1, we constructed the I-WhoQA dataset, ensuring that all examples can be answered using the model’s parametric knowledge. To further assess the trade-off between extrinsic retrieval and parametric recall, we created the I-WhoQA-Irrelevant subset. In this subset, instead of providing context relevant to the current question, we inserted entirely irrelevant content to evaluate how models handle unrelated information, and whether STRING remains effective when answering the question requires parametric knowledge rather than context-based extrinsic retrieval.
In addition to the I-WhoQA-Irrelevant subset, we construct a more realistic dataset based on HotpotQA (Yang et al. 2018) to simulate scenarios where the context appears relevant to the question but ultimately fails to support the correct answer. From the HotpotQA dataset, we derive two subsets by analyzing the relationship between the external context, the model’s parametric knowledge, and the reference answer.
First, to distinguish the role of external context from that of the model’s parametric knowledge, we ensured that the answers derived from these two sources are not the same. Building on this distinction, we obtained two final subsets. The first subset, referred to as the HotpotQA-Context subset, contains examples where the reference answer can be derived from the given contexts rather than the parametric knowledge. The second subset, known as the HotpotQA-Parametric subset, consists of examples where the reference answer matches the model’s parametric knowledge. We constructed a dataset consisting of 200 examples for each subset, except for the HotpotQA-Parametric subset for Llama3.1-8B-Instruct, which contains a total of 93 examples.
Experiment
@@
We use the same backend models and generation configuration as described in Section 1, applied to the newly constructed datasets. For STRING, we configure the local value to 128 and two different shift-ratios: 0.25 and 0.33. We report the average performance across these two settings. More details about STRING can be found in An et al. (2024b).
@@
As shown in Figure 3, STRING consistently underperforms the baseline RoPE method in terms of accuracy. This performance degradation can be attributed to STRING’s stronger extrinsic retrieval bias in long-context settings, which causes LCLMs to incorrectly focus on irrelevant context. Even when the external context is entirely unrelated to the question, STRING tends to extract the answer from the context rather than leveraging the model’s parametric knowledge, which actually contains the correct answer.